[Hallc_running] [New Logentry] RC daily update

simona at jlab.org simona at jlab.org
Fri Jul 12 09:15:01 EDT 2019

Logentry Text:
We locked up the hall yesterday towards the end of the day shift after the successful temporary fix of the HMS LCW leak. Hall C has been ready for beam since.

At the MCC this morning we were told the implementation of a the new cycling procedure for the magnets in the machine on the 20 A power supplies was a success: it's expected that future cyclings will not affect beam profile/parameters anymore and now these quads make indeed the field that's requested. The tuning of the machine has been progressing well over night and we are few hours away from getting (hopefully) high current beam in Hall C. The optics expert made the point that the machine behaves now pretty close to design, like never before since the beginning of the 12 GeV program.   

Once beam comes back we will start data taking at the last SHMS angle setting for the completion of the middle epsilon point (Q2 = 0.375 and W = 2.2). Then we will change angles and momenta for both spectrometers to complete the data taking for the low epsilon point (Q2 = 1.45 and W = 2.02). Sometime during this weekend we may request a pass change for Hall C (going from 4 to 5 pass). At that point we will make an access to allow rotation of HMS to small angle.

We will meet today at 4:30 pm on the second floor of the counting house: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3702860
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190712/d9ea5ed9/attachment.html>

More information about the Hallc_running mailing list